# The Self-Duality Equations on a Riemann Surface and Four-Dimensional Chern–Simons Theory

### v1

@@ 5 Discussion and Future Directions
In this work, we first constructed an action for the Hitchin’s equations in §2.1. We then derived this action from 4d Chern–Simons theory in §2.3. The symplectic structure arising from this four-dimensional Chern–Simons theory in §2.4 coincides with the symplectic form on the space of solutions of Hitchin’s equations that is the Kahler form on the realization of $\mathscr{M}_{\text{H}}(\Sigma,G)$ as a moduli space of Higgs bundles. We further constructed the Hitchin Hamiltonians in terms of the four-dimensional gauge field. To realize the other $\mathbb{CP}^{1}$-family of symplectic structures, we constructed corresponding four-dimensional Chern–Simons theory, each corresponding to its own family of 2d actions, with Hitchin’s equations as their equations of motion, giveing the whole hyperkähler family of symplectic structure on Hitchin moduli space. These are parametrized by $\mathbb{CP}^{1}$ which provides the base of the holomorphic fibration of the twistor space of the Hitchin moduli space. This $\mathbb{CP}^{1}$ base is thereby identified with the holomorphic plane of four-dimensional Chern–Simons theory on which the spectral parameter lives.
A similar identification between the $\mathbb{CP}^{1}$ base of the twistor space of the Hitchin moduli space and the spectral parameter Riemann sphere was made in [FrostGurdoganMason202306]. There the case of the Hitchin system arising in the Alday–Maldacena formulation of minimal surfaces in AdS was studied [AldayMaldacena200904, Alday:2009dv]. These in fact correspond to $SU(2)$ Hitchin systems with $\mathbb{Z}_{2}$ symmetry for minimal surfaces in AdS3 and $SU(4)$ with $\mathbb{Z}_{4}$ symmetry for minimal surfaces in AdS as examples of our affine Toda treatment in §4. In that problem, the the moduli space could be identified with the boundary conditions of the minimal surfaces and needs a nontrivial function $F(w)$ to appear dynamically. This work formed part of the motivation for this work, so as to find action formulations that interpolate between that of the minimal surface and the twistor description of the moduli space. Eventually one might be able to use such a description of the twistor space for the moduli space as a route to quantization of such strings in AdS. These might then have application to the study of Wilson loops and amplitudes at finite coupling.
@@
Much theory of integrable systems remains to be formulated in this 4d framework. The key 2d models of [MasonWoodhouse199605, §6] such as the Korteweg de Vries and Nonlinear Schrödinger equations have yet to be given an efficient 4d Chern–Simons formulation. Their hierarchies and those of other models are commuting flows arise from the sequences of conserved quantities. These fit naturally into the self-duality and twistor framework, for example see the descriptions in of [MasonWoodhouse199605, §8]. It seems likely that in 4d Chern–Simons theory they should arise by using the freedom the boundary conditions in (2.15): given a complex structure on $\Sigma$ with local holomorphic coordinate $w$, there is a natural hierarchy of choices of boundary condition when $\omega$ has a pole of order $q$, at $z=0$,
@@
The reductions to one dimension are also interesting, as discussed in [MasonWoodhouse199605, §7], including Nahm’s equations, spinning tops, and Painlevé equations, etc.
Ultimately, the most interesting applications of these ideas will be to the quantization of these systems, both directly, and via a quantization of the twistor formulation of their moduli spaces.

### v2

@@ 5 Discussion and Future Directions
In this work, we first constructed an action for Hitchin’s equations in §2.1. We then derived this action from 4d Chern–Simons theory in §2.3. The symplectic structure arising from this four-dimensional Chern–Simons theory in §2.4 coincides with the symplectic form on the space of solutions of Hitchin’s equations that is the Kähler form on the realization of $\mathscr{M}_{\text{H}}(\Sigma,G)$ as a moduli space of Higgs bundles. We further constructed the Hitchin Hamiltonians in terms of the four-dimensional gauge field. To realize the other $\mathbb{CP}^{1}$-family of symplectic structures, we constructed corresponding four-dimensional Chern–Simons theory, each corresponding to its own family of 2d actions, with Hitchin’s equations as their equations of motion, giving the whole hyperkähler family of symplectic structures on Hitchin moduli space. These are parametrized by $\mathbb{CP}^{1}$, which provides the base of the holomorphic fibration of the twistor space of the Hitchin moduli space. This $\mathbb{CP}^{1}$ base is thereby identified with the holomorphic plane of four-dimensional Chern–Simons theory on which the spectral parameter lives.
A similar identification between the $\mathbb{CP}^{1}$ base of the twistor space of the Hitchin moduli space and the spectral parameter Riemann sphere was made in [FrostGurdoganMason202306]. There, the case of the Hitchin system arising in the Alday–Maldacena formulation of minimal surfaces in AdS was studied [AldayMaldacena200904, AldayGaiottoMaldacena200901]. These, in fact, correspond to $\mathrm{SU}(2)$ Hitchin systems with $\mathbb{Z}_{2}$ symmetry for minimal surfaces in AdS3 and $\mathrm{SU}(4)$ with $\mathbb{Z}_{4}$ symmetry for minimal surfaces in AdS as examples of our affine Toda treatment in §4. In that problem, the moduli space could be identified with the boundary conditions of the minimal surfaces and needs a nontrivial function $F(w)$ to appear dynamically. This work formed part of the motivation for this work, so as to find action formulations that interpolate between that of the minimal surface and the twistor description of the moduli space. Eventually, one might be able to use such a description of the twistor space for the moduli space as a route to the quantization of such strings in AdS. These might then have application to the study of Wilson loops and amplitudes at finite coupling.
@@
Much of the theory of integrable systems remains to be formulated in this 4d framework. The key 2d models of [MasonWoodhouse199605, §6], such as the Korteweg de Vries and Nonlinear Schrödinger equations have yet to be given an efficient 4d Chern–Simons formulation. Their hierarchies, and those of other models, are commuting flows that arise from the sequences of conserved quantities. These fit naturally into the self-duality and twistor framework, for example, see the descriptions in of [MasonWoodhouse199605, §8]. It seems likely that in 4d Chern–Simons theory they should arise by using the freedom of the boundary conditions in (2.15): given a complex structure on $\Sigma$ with local holomorphic coordinate $w$, there is a natural hierarchy of choices of boundary condition when $\omega$ has a pole of order $q$, at $z=0$,
@@
The reductions to one dimension are also interesting, as discussed in [MasonWoodhouse199605, §7], including Nahm’s equations, spinning tops, and Painlevé equations.
Ultimately, the most interesting applications of these ideas will be to the quantization of these systems, both directly and via a quantization of the twistor formulation of their moduli spaces.
Acknowledgement.
We would like to thank Kevin Costello and David Gaiotto for comments on a draft. Furthermore, we are grateful to Vincent Caudrelier, Derek Harland, and Benoît Vicedo for many useful comments and helpful clarification regarding Remark 1.1 which appeared in an earlier version of this work. Research at Perimeter Institute is supported by the Government of Canada through Industry Canada and by the Province of Ontario through the Ministry of Research and Innovation. LJM would like to thank the Simons Collaboration on Celestial Holography CH-00001550-11 and STFC for financial support from grant number ST/T000864/1. The research of SFM is funded by the ERC Consolidator Grant #864828 “Algebraic Foundations of Supersymmetric Quantum Field Theory” (SCFTAlg).